OpenStudy (anonymous):

MeDaL will be awarded! Help!! I am given a line, y-10=2/3(x-4) How do I write a perpendicular to it line, that passes through (4,2) , in a POINT SLOPE form.

OpenStudy (anonymous):

I can find a perpendicular slope, is it -3/2?

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

i think point slope form is y = mx + b right?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

I think it is a slope intercept form, isn't it?

undeadknight26 (undeadknight26):

I honestly have no idea how to answer this question,,,@texaschic101 i need reinforcements!

undeadknight26 (undeadknight26):

@texaschic101

OpenStudy (anonymous):

@texaschic101

undeadknight26 (undeadknight26):

hehe sorry about disturbing you texas...

OpenStudy (texaschic101):

point slope form is y - y1 = m(x - x1) perpendicular line will have negative reciprocal of the given slope, therefore, the slope you need is -3/2 (4,2) x1 = 4 and y1 = 2 now we sub... y - y1 = m(x - x1) y - 2 = -3/2(x - 4) <-- point slope form

OpenStudy (anonymous):

That's it, this is the answer?

OpenStudy (texaschic101):

yes...if they want it in point slope form...now if they want it in slope intercept form, we would have to solve for y. The answer in point slope form is : y - 2 = -3/2(x - 4)

OpenStudy (anonymous):

Ty! I gave you 6-7 medals, right? Id care about medals, I just need help. Ty for helping me!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

no problem...glad to help :)